Snowfalls occur randomly and independently over the course of winter in a Minesota city. The average is one snowfall every 3 days.

a. What is the probability of five snowfalls in 2 weeks?

b. Find the probability of a snowfall today.


1
Expert's answer
2021-11-08T16:12:29-0500

Let "X=" the number of snowfalls: "X\sim Po(\lambda t)."

a.


"\lambda t=1\cdot\dfrac{14}{3}=\dfrac{14}{3}"

"P(X=5)=\dfrac{e^{-14/3}\cdot(\dfrac{14}{3})^5}{5!}\approx0.1734"



b.


"\lambda t=1\cdot\dfrac{1}{3}=\dfrac{1}{3}"

"P(X=1)=\dfrac{e^{-1/3}\cdot(\dfrac{1}{3})^1}{1!}\approx0.2388"